In a jib crane, the jib and tie rod are 5 m and 4 m long respectively. The height of crane post is 3 m and tie rod remains horizontal. Determine the forces produced in the jib and tie rod when a load of 2 kN is suspended at the crane head



1
Expert's answer
2021-10-10T09:34:02-0400


P1 and P2 are forces along the axes of the rods. Since P1 and P2 were caused by the load W, we can consider this as a triangle of forces and apply law of sines:


"\frac{P_1}{\sin 270° }= \frac{P_2}{\sin 53.13°} = \frac{2}{\sin36.87°},\\\space\\
P_1 = 2\frac{\sin 270°}{\sin 36.87°}= –3.33 \text{ kN},\\\space\\
P_2 = 2\frac{\sin 53.13°}{\sin 36.87°}= 2.67 \text{ kN}."